# Sketch2PoseNet: Efficient and Generalized Sketch to 3D Human Pose Prediction

### v1

@@ 1. Introduction
Human pose estimation holds significant importance and finds widespread application across numerous scenarios, including 3D human reconstruction (Tan et al., 2020; Zhu et al., 2016), 3D human generation (Zhuang et al., 2025; Wang et al., 2025; Zeng et al., 2023), view synthesis (Zhu et al., 2018), and animation (Zhu et al., 2024a; Liao et al., 2020). Among the various sources used for pose estimation, sketches emerge as an efficient and versatile entity. Sketches are data that can be more easily designed by artists and are widely used in animation and film production. More broadly, the term ‘sketch’ encompasses a diverse range of graphical styles, including charcoal sketches, cartoons, stick figures, kids’ drawings, oil paintings, ink paintings, and so forth.
Estimating human poses from sketches presents a significant challenge. Generalized photo-based pose estimation methods fall short in this task due to their exclusive training on realistic data. By contrast, sketches often disregard human proportionality and geometric perspective, opting for a more abstract representation of poses, thereby exacerbating the complexity of the sketch-to-pose conversion. To tackle this, Brodt et al.introduced Sketch2Pose (Brodt and Bessmeltsev, 2022), which initializes by predicting 2D joint positions from sketches and subsequently aligns a 3D parametric human model to their bones via an optimization framework. Nonetheless, this method is slow and mostly tailored towards hand-drawn sketch lines. Pursuing a swift and highly generalized solution for the sketch-to-pose task remains an open problem.
To tackle this problem, we embraced a ”learn from synthesis” strategy, which has been successfully applied in avatar modeling (Guo et al., 2023; Zhuang et al., 2024) and street view synthesis (Zhu et al., 2024c). Starting from a modest quantity of sketches and corresponding 2D human pose datasets, a large-scale sketch-3D pose dataset is synthesized by a fine-tuned image generative model conditioned on human poses. Such data synthesis is tailored for the sketch-to-pose task. Specifically, we incorporated pose perturbations to create data representing disproportionate human figures and misaligned perspectives in sketches. Furthermore, we amassed a substantial collection of sketches encompassing diverse styles, conducted detailed categorical analyses, and thereby enriched the stylistic variety of the sketches we generated. Ultimately, we produced $120,000$ such high-quality sketch-pose data pairs.
@@ 2.1. Sketch-Based Character Posing
Sketch-based character posing provides an intuitive means for users to manipulate 3D human poses, yet it introduces several significant challenges. Depth ambiguity, anatomical distortions, missing details, and diverse sketching styles make pose inference particularly difficult. Early works focused on stick figures (Hecker and Perlin, 1992; Davis et al., 2003; Mao et al., 2005; Lin et al., 2010), silhouettes (Won and Lee, 2016), and clean vector drawings (Bessmeltsev et al., 2016). These approaches, though efficient in constrained scenarios, are hindered by their reliance on unambiguous, clean inputs. For instance, Gesture3D (Bessmeltsev et al., 2016) reconstructs poses from vector drawings but assumes minimal noise, precise connectivity, and no extra strokes—requirements that are rarely met by natural, user-drawn sketches. This reliance on specific input types significantly limits the usability of such systems, as users cannot freely use diverse sketching styles to specify desired poses.
Recent approaches like Sketch2Pose (Brodt and Bessmeltsev, 2022) use a neural network to predict bitmap representations and optimize 3D model parameters for pose inference. However, due to the scarcity of sketch-to-3D model pairs for training, the method requires additional optimization to produce acceptable results. This not only introduces a significant computational burden, but also raises concerns about the reliability of the generated poses, which may lack naturalness or anatomical correctness.
An important application in this domain is the development of interactive systems for engaging with sketches. To achieve efficient inference, systems like MonsterMash (Dvorožňák et al., 2020) and Motion Doodles (Thorne et al., 2004) offer fast, intuitive sketch-based interactions but are limited by strict input formats or detailed annotations. Systems designed for articulated human poses, like those by Unlu et al.(Unlu et al., 2022) and Schmitz et al.(Schmitz et al., 2023), impose further input constraints, requiring sketches to consist of 3D primitives. Previous methods have imposed a trade-off between efficiency and input diversity due to the lack of paired sketch-3D pose datasets. In contrast, our approach addresses this gap by proposing a large-scale dataset and building a pose estimation network that directly predicts poses from sketches. This enables efficient near-real-time performance while maintaining generalizability, offering a simple, direct, and scalable solution for sketch-to-pose estimation.
@@ 2.2. Human Pose from a Single Photograph
Estimating 3D human poses from a monocular image has been extensively studied in computer vision due to its significant applications in computer graphics, animation, and human-computer interaction. Early methods relied on handcrafted features (Balan et al., 2007; Ramanan, 2011; Andriluka et al., 2010), using probabilistic models and tree-based structures. However, they struggled with occlusions, ambiguous poses, and appearance variations.
The introduction of deep learning shifted the field, with DeepPose (Toshev and Szegedy, 2014) being one of the first CNN-based approaches. This was followed by methods like Tekin et al. (Tekin et al., 2016), which integrated CNNs with structured prediction to improve pose accuracy, and Martinez et al. (Martinez et al., 2017), which proposed a fully connected network for 2D-to-3D lifting. Zhou et al. (Zhou et al., 2017) added geometric constraints, while Pavlakos et al. (Pavlakos et al., 2017) used volumetric heatmaps for joint localization.
Despite progress, the need for large labeled datasets limited generalization, especially for non-photorealistic inputs. The introduction of parametric models like SMPL (Bogo et al., 2016) and SMPL-X (Pavlakos et al., 2019) advanced pose and shape estimation with 3D human priors. Many methods (Zhang et al., 2021; Li et al., 2021, 2022; Goel et al., 2023) focus on improving the accuracy of human mesh recovery. Weakly supervised approaches like HMR (Kanazawa et al., 2018) regressed SMPL parameters using 2D keypoints and adversarial losses, and HMD (Zhu et al., 2021, 2019) further refined detailed shape based on the predicted SMPL mesh. Kolotouros et al.’s SPIN (Kolotouros et al., 2019) refined this approach with optimization, while EFT (Joo et al., 2021) fine-tuned SMPL predictions. Methods like 3DCrowdNet (Choi et al., 2022), JOTR (Li et al., 2023b), and DPMesh (Zhu et al., 2024b) are designed to recover the occluded body mesh. Self-supervised methods, such as Wang et al. (Wang et al., 2019) and Novotny et al. (Novotny et al., 2019), reduced dependence on labeled data using geometric consistency.
Our method bridges human pose estimation and character drawing by leveraging visual priors in pre-trained diffusion models, as seen in VPD (Zhao et al., 2023). Using our dataset, our fine-tuned network extracts structural and spatial features from the denoising U-Net for accurate human mesh recovery. By processing the input image in a single inference pass (Zhu et al., 2024b), our approach adapts diffusion priors to handle abstract sketches, ensuring reliable pose estimation.
@@ 3. SKEP-120K Dataset
It is widely recognized that the quality and abundance of the training data heavily influence the success of learning-based techniques for human mesh recovery. Surprisingly, we find that there is currently a notable absence of a large-scale, high-quality dataset containing sketches and 3D human poses. Though several available datasets (Madhu et al., 2022; Brodt and Bessmeltsev, 2022; Smith et al., 2023; Ju et al., 2023) offer a substantial number of sketches, they provide only 2D pose labels and typically a single sketch style, making them insufficient for training a highly accurate, generalizable sketch-to-pose model.
@@
Therefore, we propose a Sketch and 3D Pose dataset with $120k$ data pairs in various sketch styles, named as SKEP-120K dataset. As shown in Fig. 3, the dataset encompasses six styles according to artificial human scenes: cartoons, oil paintings, ink paintings, charcoal sketches, stick figures, and kids’ drawings. Each style contains approximately 20,000 images. Our dataset provides human bounding boxes, 16 human joints (both 2D and 3D, with corresponding visible/invisible/included attributes), SMPL pose parameters, and text information. Due to the different definitions of the skeleton for a human pose in different datasets and considering the characteristics of gesture expression in the sketch, we define a new 3D skeleton to represent the human pose. Specifically, the body parts are based on MSCOCO database (Lin et al., 2014), and two additional joints regarding left and right toes are added to reflect fine-scale leg poses.
@@
The creation process of our dataset is shown in Fig. 2. Firstly, we utilize VPoser (Pavlakos et al., 2019) to generate random SMPL models exhibiting diverse and plausible poses, where a variational autoencoder is designed to capture latent representations of human poses. Given the pervasive foreshortening in sketches, which makes the depicted human structure to deviate from the standard 3D-to-2D projection observed in real-captured images, we introduce skeletal-proportion perturbations during the 3D-to-2D mapping by adding random biases to the projected limb lengths as a skeleton-level data augmentation. This strategy yields skeletal configurations that better capture the characteristic proportional exaggerations and imbalances of hand-drawn sketches, thereby enhancing the pose diversity and accuracy of the sketch dataset and producing 2D joint annotations that are better aligned with sketch scenarios. Next, we consolidate data from the Sketch2Pose  (Brodt and Bessmeltsev, 2022), Human-Art  (Ju et al., 2023), and Amateur Drawing datasets  (Smith et al., 2023), unify their 2D keypoint annotations to our defined set of 16 joints, and partition them into six groups by drawing style. We then leverage BLIP2 (Li et al., 2023a) to generate appearance descriptions on sketch images and motion descriptions on SMPL rendering images. These descriptions, together with the target style labels, serve as conditioning prompts, yielding a sketch training set. On this basis, we train a text-conditioned image generation model following ControlNet (Zhang et al., 2023) to synthesize sketch data. During training, sketches from all six styles and their associated annotation parameters are jointly used to train the same model.
@@
After that, we manually curate the model-generated dataset to improve its quality further. Specifically, we invite experienced 3D modelers to filter out approximately $10\%$ of the sketches in each style, including those with severely cluttered backgrounds that compromise subject visibility, those showing pose inconsistencies caused by overly challenging conditioning poses, and a subset with extremely distorted poses introduced by the added random biases. As a result, we generate character images across various styles with high accuracy and adherence to the 2D skeleton distribution. Given the diverse line-based nature of sketches, we use the off-the-shelf outline detector (Canny, 1986) to extract line distributions within these images. A threshold is applied to identify the smallest region encompassing most lines, defining the character’s bounding box. We compare the human bounding box with the bounding box generated from the detector of Human-Art, leaving a more accurate result, and manually filtering out the undetectable cases. We also detect the occlusion of each joint based on the occlusion relationship of the SMPL mesh, which is recorded as the label of each joint.
@@ 4. Method
Drawing inspiration from VPD (Zhao et al., 2023), our core idea involves extracting high-level pre-trained knowledge from a diffusion model. A fundamental prerequisite for achieving this is the extraction of 2D guidance.
The first step in extracting 2D guidance is to estimate 2D joints from input sketches. Leveraging our proposed SKEP-120K dataset, we have fine-tuned two state-of-the-art network models for human detection and 2D joint extraction from sketches. Specifically, the input sketches are first resized and padded to the resolution of $256\times 192$ pixels to preserve the aspect ratio. Then, the human detector YOLOX (Redmon, 2016) is fine-tuned by Human-Art dataset for human’s bounding box detection in sketches, and the ViTPose (Xu et al., 2022) model is fine-tuned for 2D joints prediction from these bounded sketches. ViTPose utilizes a straightforward, non-hierarchical vision transformer as the encoder to capture human features in drawings, combined with a lightweight decoder that predicts body joints in a top-down approach. Finally, we obtain 2D joints $J^{2D}\in\mathbb{R}^{K\times 2}$ along with their corresponding confidence and transform them into heatmaps $H_{j}^{2D}\in\mathbb{R}^{K\times H^{{}^{\prime}}\times W^{{}^{\prime}}}$ using 2D Gaussian kernels (Cheng et al., 2020).
After the 2D joints are obtained, pose features are extracted from the 2D joints $J^{2D}$ and heatmaps $H_{j}^{2D}$, which provides spatial guidance for the denoising U-Net (Rombach et al., 2022) backbone $\boldsymbol{\epsilon_{\theta}}$. This process is referred as $p_{\phi_{1}}(\mathcal{G}\!\mid\!\boldsymbol{x})$ in Eq. 1. For the input image $\boldsymbol{x}$, we convert the cropped image $\boldsymbol{x}\in\mathbb{R}^{H\times W\times C}$ from pixel space to the latent space with frozen encoder $\mathcal{E}$ in the trained VQGAN from the Controlnet framework to obtain the latent representation $z_{0}\in\mathbb{R}^{H^{{}^{\prime}}\times W^{{}^{\prime}}\times\mathcal{G}_{z}}$. Then, we concatenate the heatmap $H_{j}^{2D}$ with the input image $z_{0}$ to obtain $\mathcal{G}_{f}\in\mathbb{R}^{(K+{\mathcal{G}_{z}})\times H^{{}^{\prime}}\times W^{{}^{\prime}}}$. In most previous diffusion models  (Zhang et al., 2023; Rombach et al., 2022; Dhariwal and Nichol, 2021), the prompt guidance $\mathcal{G}_{t}$ usually relies on text embeddings derived from a frozen CLIP (Radford et al., 2021) model. In contrast, we replace the text with 2D joint positions $J^{2D}$ as verified in  (Zhu et al., 2024b). To match the text token dimension $D_{j}$, a two-layer MLP is used to enhance the dimensionality of the 2D joint positions to 768 in the pre-trained diffusion model. This generates a spatial guidance $\mathcal{G}_{t}\in\mathbb{R}^{K\times D_{j}}$. The process can be expressed as follows:

### v2

@@ 1. Introduction
Human pose estimation holds significant importance and finds widespread application across numerous scenarios, including 3D human reconstruction (tan2020self; zhu2016video), 3D human generation (zhuang2025idol; wang2025tera; zeng2023avatarbooth), view synthesis (zhu2018view), and animation (zhu2024champ; liao2020speech2video). Among the various sources used for pose estimation, sketches emerge as an efficient and versatile entity. Sketches are data that can be more easily designed by artists and are widely used in animation and film production. More broadly, the term ‘sketch’ encompasses a diverse range of graphical styles, including charcoal sketches, cartoons, stick figures, kids’ drawings, oil paintings, ink paintings, and so forth.
Estimating human poses from sketches presents a significant challenge. Generalized photo-based pose estimation methods fall short in this task due to their exclusive training on realistic data. By contrast, sketches often disregard human proportionality and geometric perspective, opting for a more abstract representation of poses, thereby exacerbating the complexity of the sketch-to-pose conversion. To tackle this, Brodt et al.introduced Sketch2Pose (brodt2022sketch2pose), which initializes by predicting 2D joint positions from sketches and subsequently aligns a 3D parametric human model to their bones via an optimization framework. Nonetheless, this method is slow and mostly tailored towards hand-drawn sketch lines. Pursuing a swift and highly generalized solution for the sketch-to-pose task remains an open problem.
To tackle this problem, we embraced a ”learn from synthesis” strategy, which has been successfully applied in avatar modeling (guo2023rafare; zhuang2024towards) and street view synthesis (zhu2024streetsyn). Starting from a modest quantity of sketches and corresponding 2D human pose datasets, a large-scale sketch-3D pose dataset is synthesized by a fine-tuned image generative model conditioned on human poses. Such data synthesis is tailored for the sketch-to-pose task. Specifically, we incorporated pose perturbations to create data representing disproportionate human figures and misaligned perspectives in sketches. Furthermore, we amassed a substantial collection of sketches encompassing diverse styles, conducted detailed categorical analyses, and thereby enriched the stylistic variety of the sketches we generated. Ultimately, we produced $120,000$ such high-quality sketch-pose data pairs.
@@ 2.1. Sketch-Based Character Posing
Sketch-based character posing provides an intuitive means for users to manipulate 3D human poses, yet it introduces several significant challenges. Depth ambiguity, anatomical distortions, missing details, and diverse sketching styles make pose inference particularly difficult. Early works focused on stick figures (Hecker1992; Davis2003; Mao2005a; Lin2010), silhouettes (shadowTheatre), and clean vector drawings (Bessm2016:10.1145/2980179.2980240). These approaches, though efficient in constrained scenarios, are hindered by their reliance on unambiguous, clean inputs. For instance, Gesture3D (Bessm2016:10.1145/2980179.2980240) reconstructs poses from vector drawings but assumes minimal noise, precise connectivity, and no extra strokes—requirements that are rarely met by natural, user-drawn sketches. This reliance on specific input types significantly limits the usability of such systems, as users cannot freely use diverse sketching styles to specify desired poses.
Recent approaches like Sketch2Pose (brodt2022sketch2pose) use a neural network to predict bitmap representations and optimize 3D model parameters for pose inference. However, due to the scarcity of sketch-to-3D model pairs for training, the method requires additional optimization to produce acceptable results. This not only introduces a significant computational burden, but also raises concerns about the reliability of the generated poses, which may lack naturalness or anatomical correctness.
An important application in this domain is the development of interactive systems for engaging with sketches. To achieve efficient inference, systems like MonsterMash (dvorovzvnak2020monster) and Motion Doodles (thorne2004motion) offer fast, intuitive sketch-based interactions but are limited by strict input formats or detailed annotations. Systems designed for articulated human poses, like those by Unlu et al.(unlu2022interactive) and Schmitz et al.(schmitz2023interactive), impose further input constraints, requiring sketches to consist of 3D primitives. Previous methods have imposed a trade-off between efficiency and input diversity due to the lack of paired sketch-3D pose datasets. In contrast, our approach addresses this gap by proposing a large-scale dataset and building a pose estimation network that directly predicts poses from sketches. This enables efficient near-real-time performance while maintaining generalizability, offering a simple, direct, and scalable solution for sketch-to-pose estimation.
@@ 2.2. Human Pose from a Single Photograph
Estimating 3D human poses from a monocular image has been extensively studied in computer vision due to its significant applications in computer graphics, animation, and human-computer interaction. Early methods relied on handcrafted features (balan2007detailed; Ramanan2011; andriluka2010monocular), using probabilistic models and tree-based structures. However, they struggled with occlusions, ambiguous poses, and appearance variations.
The introduction of deep learning shifted the field, with DeepPose (toshev2014deeppose) being one of the first CNN-based approaches. This was followed by methods like Tekin et al. (Tekin_2016_CVPR), which integrated CNNs with structured prediction to improve pose accuracy, and Martinez et al. (martinez_2017_3dbaseline), which proposed a fully connected network for 2D-to-3D lifting. Zhou et al. (Zhou_2017_ICCV) added geometric constraints, while Pavlakos et al. (pavlakos2017coarse) used volumetric heatmaps for joint localization.
Despite progress, the need for large labeled datasets limited generalization, especially for non-photorealistic inputs. The introduction of parametric models like SMPL (bogo2016keep) and SMPL-X (pavlakos2019expressive) advanced pose and shape estimation with 3D human priors. Many methods (zhang2021pymaf; li2021hybrik; li2022cliff; goel2023humans) focus on improving the accuracy of human mesh recovery. Weakly supervised approaches like HMR (KanazawaBJM18) regressed SMPL parameters using 2D keypoints and adversarial losses, and HMD (zhu2021detailed; zhu2019detailed) further refined detailed shape based on the predicted SMPL mesh. Kolotouros et al.’s SPIN (kolotouros2019spin) refined this approach with optimization, while EFT (joo2020eft) fine-tuned SMPL predictions. Methods like 3DCrowdNet (choi2022learning), JOTR (li2023jotr), and DPMesh (zhu2024dpmesh) are designed to recover the occluded body mesh. Self-supervised methods, such as Wang et al. (wang20193d) and Novotny et al. (Novotny2019), reduced dependence on labeled data using geometric consistency.
Our method bridges human pose estimation and character drawing by leveraging visual priors in pre-trained diffusion models, as seen in VPD (zhao2023unleashing). Using our dataset, our fine-tuned network extracts structural and spatial features from the denoising U-Net for accurate human mesh recovery. By processing the input image in a single inference pass (zhu2024dpmesh), our approach adapts diffusion priors to handle abstract sketches, ensuring reliable pose estimation.
@@ 3. SKEP-120K Dataset
It is widely recognized that the quality and abundance of the training data heavily influence the success of learning-based techniques for human mesh recovery. Surprisingly, we find that there is currently a notable absence of a large-scale, high-quality dataset containing sketches and 3D human poses. Though several available datasets (madhu2022enhancing; brodt2022sketch2pose; smith2023method; ju2023human) offer a substantial number of sketches, they provide only 2D pose labels and typically a single sketch style, making them insufficient for training a highly accurate, generalizable sketch-to-pose model.
@@
Therefore, we propose a Sketch and 3D Pose dataset with $120k$ data pairs in various sketch styles, named as SKEP-120K dataset. As shown in Fig. 3, the dataset encompasses six styles according to artificial human scenes: cartoons, oil paintings, ink paintings, charcoal sketches, stick figures, and kids’ drawings. Each style contains approximately 20,000 images. Our dataset provides human bounding boxes, 16 human joints (both 2D and 3D, with corresponding visible/invisible/included attributes), SMPL pose parameters, and text information. Due to the different definitions of the skeleton for a human pose in different datasets and considering the characteristics of gesture expression in the sketch, we define a new 3D skeleton to represent the human pose. Specifically, the body parts are based on MSCOCO database (lin2014microsoft), and two additional joints regarding left and right toes are added to reflect fine-scale leg poses.
@@
The creation process of our dataset is shown in Fig. 2. Firstly, we utilize VPoser (pavlakos2019expressive) to generate random SMPL models exhibiting diverse and plausible poses, where a variational autoencoder is designed to capture latent representations of human poses. Given the pervasive foreshortening in sketches, which makes the depicted human structure to deviate from the standard 3D-to-2D projection observed in real-captured images, we introduce skeletal-proportion perturbations during the 3D-to-2D mapping by adding random biases to the projected limb lengths as a skeleton-level data augmentation. This strategy yields skeletal configurations that better capture the characteristic proportional exaggerations and imbalances of hand-drawn sketches, thereby enhancing the pose diversity and accuracy of the sketch dataset and producing 2D joint annotations that are better aligned with sketch scenarios. Next, we consolidate data from the Sketch2Pose  (brodt2022sketch2pose), Human-Art  (ju2023human), and Amateur Drawing datasets  (smith2023method), unify their 2D keypoint annotations to our defined set of 16 joints, and partition them into six groups by drawing style. We then leverage BLIP2 (li2023blip) to generate appearance descriptions on sketch images and motion descriptions on SMPL rendering images. These descriptions, together with the target style labels, serve as conditioning prompts, yielding a sketch training set. On this basis, we train a text-conditioned image generation model following ControlNet (zhang2023adding) to synthesize sketch data. During training, sketches from all six styles and their associated annotation parameters are jointly used to train the same model.
@@
After that, we manually curate the model-generated dataset to improve its quality further. Specifically, we invite experienced 3D modelers to filter out approximately $10\%$ of the sketches in each style, including those with severely cluttered backgrounds that compromise subject visibility, those showing pose inconsistencies caused by overly challenging conditioning poses, and a subset with extremely distorted poses introduced by the added random biases. As a result, we generate character images across various styles with high accuracy and adherence to the 2D skeleton distribution. Given the diverse line-based nature of sketches, we use the off-the-shelf outline detector (canny1986computational) to extract line distributions within these images. A threshold is applied to identify the smallest region encompassing most lines, defining the character’s bounding box. We compare the human bounding box with the bounding box generated from the detector of Human-Art, leaving a more accurate result, and manually filtering out the undetectable cases. We also detect the occlusion of each joint based on the occlusion relationship of the SMPL mesh, which is recorded as the label of each joint.
@@ 4. Method
Drawing inspiration from VPD (zhao2023unleashing), our core idea involves extracting high-level pre-trained knowledge from a diffusion model. A fundamental prerequisite for achieving this is the extraction of 2D guidance.
The first step in extracting 2D guidance is to estimate 2D joints from input sketches. Leveraging our proposed SKEP-120K dataset, we have fine-tuned two state-of-the-art network models for human detection and 2D joint extraction from sketches. Specifically, the input sketches are first resized and padded to the resolution of $256\times 192$ pixels to preserve the aspect ratio. Then, the human detector YOLOX (redmon2016you) is fine-tuned by Human-Art dataset for human’s bounding box detection in sketches, and the ViTPose (xu2022vitpose) model is fine-tuned for 2D joints prediction from these bounded sketches. ViTPose utilizes a straightforward, non-hierarchical vision transformer as the encoder to capture human features in drawings, combined with a lightweight decoder that predicts body joints in a top-down approach. Finally, we obtain 2D joints $J^{2D}\in\mathbb{R}^{K\times 2}$ along with their corresponding confidence and transform them into heatmaps $H_{j}^{2D}\in\mathbb{R}^{K\times H^{{}^{\prime}}\times W^{{}^{\prime}}}$ using 2D Gaussian kernels (cheng2020higherhrnet).
After the 2D joints are obtained, pose features are extracted from the 2D joints $J^{2D}$ and heatmaps $H_{j}^{2D}$, which provides spatial guidance for the denoising U-Net (rombach2022high) backbone $\boldsymbol{\epsilon_{\theta}}$. This process is referred as $p_{\phi_{1}}(\mathcal{G}\!\mid\!\boldsymbol{x})$ in Eq. 1. For the input image $\boldsymbol{x}$, we convert the cropped image $\boldsymbol{x}\in\mathbb{R}^{H\times W\times C}$ from pixel space to the latent space with frozen encoder $\mathcal{E}$ in the trained VQGAN from the Controlnet framework to obtain the latent representation $z_{0}\in\mathbb{R}^{H^{{}^{\prime}}\times W^{{}^{\prime}}\times\mathcal{G}_{z}}$. Then, we concatenate the heatmap $H_{j}^{2D}$ with the input image $z_{0}$ to obtain $\mathcal{G}_{f}\in\mathbb{R}^{(K+{\mathcal{G}_{z}})\times H^{{}^{\prime}}\times W^{{}^{\prime}}}$. In most previous diffusion models  (zhang2023adding; rombach2022high; dhariwal2021diffusion), the prompt guidance $\mathcal{G}_{t}$ usually relies on text embeddings derived from a frozen CLIP (radford2021learning) model. In contrast, we replace the text with 2D joint positions $J^{2D}$ as verified in  (zhu2024dpmesh). To match the text token dimension $D_{j}$, a two-layer MLP is used to enhance the dimensionality of the 2D joint positions to 768 in the pre-trained diffusion model. This generates a spatial guidance $\mathcal{G}_{t}\in\mathbb{R}^{K\times D_{j}}$. The process can be expressed as follows: